A car travels 240 miles on 15 gallons of gas. How far can the car travel on 20 gallons of gas?
Answer by ptaylor(2198)


Let x=distance traveled on 20 gal
One approach is to set up a simple ratio: 15 is to 240 as 20 is to x and this is written as follows:
15/240=20/x cross multiply (actually we are multiplying both sides by 240x)
15x=240*20=4800 divide both sides by 15
x=320 mi------------------------------------distance travelled on 20 gal
Another approach would be to determine the mileage the car gets in miles/gal:
If the car goes 240 mi on 15 gal, then it goes 240/15 or 16 mi on 1 gal.
Soooooooooo our car gets 16 mi/gal and it can go 16*20 or 320 mi on 20 gal

Hope this helps---ptaylor